# Towards SLO-Optimized LLM Serving via Automatic Inference Engine Tuning

### v1

@@ Abstract
A service-level objective (SLO) is a target performance metric of service that cloud vendors aim to ensure. Delivering optimized SLOs can enhance user satisfaction and improve the competitiveness of cloud vendors. As large language models (LLMs) are gaining increasing popularity across various fields, it is of great significance to optimize SLOs for LLM inference services. In this paper, we observe that adjusting the parameters of LLM inference engines can improve service performance, and the optimal parameter configurations of different services are different. Therefore, we propose SCOOT, an automatic performance tuning system to optimize SLOs for each LLM inference service by tuning the parameters of the inference engine. We first propose a generalized formulation of the tuning problem to handle various objectives and constraints between parameters, and SCOOT exploits the Bayesian optimization (BO) technique to resolve the problem via exploration and exploitation. Moreover, SCOOT adopts a random forest to learn hidden constraints during the tuning process to mitigate invalid exploration. To improve the tuning efficiency, SCOOT utilizes the parallel suggestion to accelerate the tuning process. Extensive experiments demonstrate that SCOOT can significantly outperform existing tuning techniques in SLO optimization while greatly improving the tuning efficiency.
Index Terms:
large language model, inference engine, service-level object, performance tuning, machine learning
I Introduction
@@
With the impressive capabilities demonstrated by large language models (LLMs) across various applications [1], cloud vendors such as Alibaba Cloud and AWS have started offering services of LLM deployment and inference [2, 3]. Customers (i.e., application developers and providers) can deploy specified LLMs on their exclusive computing resources. Through the application programming interface (API) provided by the cloud vendor, customers‘ requests can be continuously served by the deployed LLMs.
@@
These LLM inference services run LLM instances with advanced inference engines such as vLLM [4] and TensorRT-LLM [5], which are equipped with cutting-edge technologies, such as continuous batching [6], paged attention [7], and chunked prefill [8]. These techniques can accelerate inference and improve throughput, thus delivering a high service performance to customers. To provide flexibility of use, inference engines expose many parameters, of which the values can be modified to enable a particular technique or to change strategies of request scheduling and memory allocation.
@@
To ensure service performance, customers always agree on service-level objectives (SLOs) with cloud vendors. SLOs are defined as a series of performance metric constraints, such as requiring that the 95th percentile latency of requests be less than 1 second. A tight SLO can gain a reputation for cloud vendors and attract more customers. Therefore, it is significant for cloud vendors to optimize SLOs (e.g., guarantee a lower tail latency) to improve customer satisfaction while enhancing market competitiveness.
@@
Existing LLM deployment solutions usually employ the default parameter configuration or adopt an empirically good configuration of the LLM inference engine for all inference services. However, in this paper, we observe that adjusting the parameters of LLM inference engines has great potential to improve service performance, and the optimal parameter configurations of various services are different. Therefore, we stand in the shoes of cloud vendors and optimize SLOs for each LLM inference service by tuning the parameters of the LLM inference engine.
@@
Numerous performance tuning methods have been proposed and applied across various fields [9], but they all fall short of efficiency and optimality for tuning LLM inference engines. Methods like random sampling and meta-heuristic algorithms, such as Monte Carlo sampling [10] and genetic algorithms [11] lack efficiency as they fail to fully utilize historical information. Besides, heuristic searches often rely on expert knowledge and elaborate pre-profiling to model relationships between parameters and performance. Nevertheless, with the rapid evolution of technologies in LLM inference engines, the parameters’ numbers and ranges are frequently updated, which makes the modeled parameter-performance relationship outdated and the designed heuristic search methods inapplicable. Learning-based approaches such as reinforcement learning (RL) [12] and Bayesian optimization (BO) [13] have also been widely exploited in performance tuning. They can effectively leverage historical information and tune parameters automatically without prior knowledge. However, existing learning-based tuning methods have difficulties tuning LLM inference engines due to the following three unique challenges.
@@
Challenge 1: Customers seek to enhance different performance metrics depending on their applications, such as improving request throughput for the offline recommendation application, reducing request tail latency for the online classification application, and minimizing time-to-first-token (TTFT) and time-per-output-token (TPOT) simultaneously for interactive applications such as chatbot. Challenge 2: The tuning process is complicated by both known constraints that can be provided to the tuner ahead of time and hidden constraints that are initially unknown and must be learned throughout the tuning process. Challenge 3: Tuning methods always require significant time to evaluate different parameter configurations, making the tuning process for a large number of LLM inference services time-prohibitive.
To address the above challenges, we propose SCOOT, a ServiCe-level Objective Oriented performance Tuning system, which automatically tune parameters of LLM inference engines to optimize SLOs for LLM inference services. We first propose a general formulation of the inference engine tuning problem to accommodate various optimization objectives and complex constraints, and SCOOT can resolve the problem with BO, where single-objective Bayesian optimization (SOBO) and multi-objective Bayesian optimization (MOBO) [14, 15, 16] are respectively employed to search optimized parameter configurations in single-objective and multi-objective situations, thus addressing challenge 1. Moreover, SCOOT prunes the search space with known constraints and exploits a random forest to learn hidden constraints during the tuning process, thus mitigating challenge 2. To resolve challenge 3, the parallel suggestion is implemented on top of BO to recommend multiple parameter configurations for evaluation each time, thus fully utilizing idle computing resources to greatly accelerate the tuning process and improve the tuning efficiency.
We conduct extensive experiments with various LLMs and different types and numbers of GPUs under request traces collected from various applications in the production environment of Ant Group. The results show that SCOOT can speed up the tuning process and significantly optimize SLOs, improving the request throughput by up to 68.3%, reducing the request tail latency by up to 40.6%, and reducing the TTFT and TPOT by up to 99.8% and 61.0%, respectively, compared to the default parameter configuration and existing tuning methods. The main contribution of this paper is listed as follows.
  • Random forest regression is employed by SCOOT to learn hidden constraints during the tuning process to avoid invalid explorations while the parallel suggestion technique is adopted to significantly improve the tuning efficiency using additional computing resources.
II Background and Motivation
II-A Research Background
TABLE I: PARAMETERS TO TUNE.
@@ II-A Research Background
II-A1 LLM Inference Engine and Parameters
Inference engines integrate numerous advanced LLM inference techniques such as continuous batching [6], paged attention [7], and chunked prefill [8] and expose many parameters. For vLLM, these parameters include boolean variables such as enable-chunked-prefill that can enable the technique of chunked prefill, integer and float variables such as max-num-seqs and scheduler-delay-factor that can change the request scheduling strategy of continuous batching, and enumeration variables such as block-size that can change the memory allocation policy of paged attention. In this paper, we focus on tuning parameters that do not affect the accuracy of LLMs. Therefore, we do not consider parameters related to model compression such as quantization. Additionally, although speculative decoding has been theoretically proven not to hurt LLM accuracy [17], it still affects the LLMs’ generation results, making it inapplicable for certain applications. Hence, we also do not tune parameters related to speculative decoding. In this paper, we choose vLLM as the inference engine, and the parameters to be tuned are listed in Table I, which constructs a huge search space of billions of configuration points.
II-A2 Service-level Objective and Stress Testing
To ensure service performance, customers establish SLOs with cloud vendors. If SLOs are violated, cloud vendors not only have to compensate customers but also face reputational damage. Additionally, customers also need to make a commitment to the request rate, guaranteeing that the request rate for the inference service will not exceed an agreed threshold. Typically, cloud vendors stress test services with a high request rate (i.e., a request rate that is close to the agreed threshold) and set the performance metrics achieved under the stress test as SLOs. This ensures that SLOs won’t be violated even under heavy workload scenarios. By optimizing the service performance under stress testing, cloud vendors can deliver better SLOs to customers.
TABLE II: VARIOUS SERVICE CHARACTERISTICS.
@@ II-B Experimental Observations
Figure 1: Optimal TTFT and TPOT for various services. The TTFT and TPOT shown are relative values compared to those of the default parameter configuration. The lower, the better.
II-B1 Parameter adjustment of inference engine can enhance performance for LLM inference services
@@
We conduct experiments to find the optimal parameter configurations of vLLM with a grid search for five services of various LLMs deployed on different numbers and types of GPUs under request traces of various applications. The service characteristics in the experiments are shown in Table II, where A10 and A100 separately represent the NVIDIA A10 24G GPU and NVIDIA A100 80G GPU, and SQL and BOT represent the request traces of text-to-sql and chatbot applications in Ant Group. Other experimental settings are the same as those in Section V. Fig.1 presents the relative TTFT and TPOT performed by optimal parameter configurations for these services. We can observe that TTFT and TPOT can be reduced by up to 98.9% and 49.9% compared to the default configuration, respectively, which confirms the significance of performance tuning.
II-B2 Different inference services’ optimal parameter configurations are different
@@
We conduct experiments to apply the optimal parameter configuration of a service to other services. Figure 2 presents the experimental results. We can observe that the performance exhibited by the optimal parameter configuration for a service may perform poorly or even cause errors for other services. Hence, there is no one “best practice” configuration that works best in all scenarios, and it is essential to conduct performance tuning for each inference service.
II-C Challenges
II-C1 Various Optimization Objectives
Customers of LLM inference services intend to employ different performance metrics as objectives for SLO optimization according to their specific application requirements. For non-interactive applications, periodically invoked offline applications such as offline recommendation, prioritize improving request throughput, while online applications such as online classification, aim to reduce request tail latency. Interactive applications such as chatbot continuously display new tokens generated by the LLM to application users, and hence they tend to simultaneously minimize two token-level metrics, time-to-first-token (TTFT), and time-per-output-token (TPOT). TTFT is the time from the user sending a request to receiving the first token, and TPOT is the average time interval between tokens. By minimizing TTFT and TPOT, the output progress can be more fluent, thereby enhancing the user experience. Therefore, optimizing SLOs for non-interactive applications often poses a single-objective optimization problem, whereas optimizing SLOs for interactive applications is a multi-objective problem.
II-C2 Complex Known and Hidden Constraints
Some parameters of LLM engines depend on the settings of other parameters, thus causing constraints on the search space. For example, for vLLM, max-num-batched-tokens must be greater than or equal to max-num-seqs. If these constraints are violated, the inference engine can not be started. We refer to these as known constraints, which can be provided to the tuner ahead of time. In addition, given a specific service, certain parameter combinations can lead to inference engine crashes during the stress testing. In this paper, these infeasible parameter combinations are referred to as hidden constraints. Different services have different hidden constraints that are initially unknown and must be learned throughout the tuning process. Specifically, for some services, vLLM often crushes due to a timeout error during the stress testing when scheduler-delay-factor is set to a large value.
(a) Relative time-to-first-token (TTFT) compared to default configuration. The lower, the better.
(b) Relative time-per-output-token (TPOT) compared to default configuration. The lower, the better.
Figure 2: TTFT and TPOT for applying optimal parameter configurations of different services to other services. Conf. $X$ represents the optimal parameter configuration for the service $X\in\{A,B,C,D,E\}$. The red lines (1.0×) in Sub-figures (a) and (b) indicate the TTFT and TPOT under the default parameter configuration of the inference engine, respectively.
II-C3 High Evaluation Overhead
Tuning methods always necessitate evaluating various parameter configurations. For the inference engine tuning, each evaluation (i.e., stress testing) can take several minutes. For example, it takes about 6 minutes to stress test a service with the LLAMA2-7B LLLM [18] deployed on 2 A10 GPUs using 1,000 requests of the BOT application. For more requests and larger LLMs, the time required for stress testing will become longer. Hence, even with only 30 evaluations, the time spent tuning a single service can reach several hours, making the time required for tuning a multitude of newly deployed services prohibitive.
II-D Limitation of Existing Solutions
II-D1 Random Sampling and Meta-Heuristics
Random sampling randomly samples points from the search space to evaluate while meta-heuristics such as genetic algorithms [11] and particle swarm optimization [15] first sample several points, and then optimize around these initially sampled points. While these methods can offer sufficient diversity in solutions, they fail to fully utilize the historical information and often require a large number of evaluations for sampled configuration points to find the optimal solution. However, for tuning the inference engine, evaluations involve stress-testing the service, leading to a high time overhead. Therefore, random sampling and meta-heuristics lack efficiency for tuning the inference engine.
II-D2 Heuristic Search Algorithms
Heuristic search algorithms always leverage rules to guide the tuning process, where the search rules are often designed according to professional domain knowledge and elaborate pre-profiling to model the effects of each parameter on performance metrics. However, with the repaid development of LLM inference techniques, inference engines update frequently. For example, vLLM updates almost every few days or weeks, which leads to updates in the number and ranges of parameters, as well as changes in the effect of parameters on performance metrics. Therefore, the rules are outdated frequently, requiring substantial manpower to develop new rules. Consequently, heuristic algorithms are not suitable for tuning the LLM inference engine.
II-D3 Existing Learning-based Approaches
RL [12] and BO [13] can effectively utilize historical information to intelligently find optimal parameter configurations without prior knowledge and additional manpower. However, RL requires a time-consuming training process to obtain a well-behaved tuning agent while existing BO solutions fail to address the three unique challenges mentioned before.
TABLE III: NOTATIONS OF PROBLEM FORMULATION.
Notation Description
$\boldsymbol{\Lambda}$ The search space of parameter configurations
$\boldsymbol{x}$ A vector denotes a specific parameter configuration
$\boldsymbol{x}[i]$ The specific value for the $i$th parameter of $\boldsymbol{x}$.
$T(\boldsymbol{x})$ The objective function of request throughput for $\boldsymbol{x}$
$L(\boldsymbol{x})$ The objective function of tail request latency for $\boldsymbol{x}$
$\Phi(\boldsymbol{x})$ The objective function of average TTFT for $\boldsymbol{x}$
$\Theta(\boldsymbol{x})$ The objective function of average TPOT for $\boldsymbol{x}$
$\lambda_{t},\lambda_{l},\lambda_{\theta},\lambda_{\phi}$ Binary variables to control optimization objectives
$\mathcal{C}$ The set of known constraints
$POF(\boldsymbol{x})$ The probability of feasibility (POF) for $\boldsymbol{x}$
$\Delta$ POF threshold to mitigate hidden constraint violations
III Problem Formulation
@@
Given an inference service, the goal of performance tuning for the LLM inference engine is to find the optimal or near-optimal parameter configuration that maximizes the objective and satisfies known and hidden constraints. As we mentioned before, different customers want to optimize different performance metrics. Therefore, we provide a generalized problem formulation that supports various optimization objectives and complex constraints. The main notations utilized in the formulation are listed in Table III.
@@
Figure 3: SCOOT workflow. SCOOT leverages BO to find optimized parameter configurations via exploration and exploitation.
@@
We define the configuration space $\boldsymbol{\Lambda}=\Lambda_{1}\times\Lambda_{2}\times...\times\Lambda_{n}$, where $\Lambda_{i}$ represent the range of the $i$th parameter to tune. Vector $\boldsymbol{x}\in\boldsymbol{\Lambda}$ is utilized to denote a specific parameter configuration, where the $i$th element $\boldsymbol{x}[i]\in\Lambda_{i}$ corresponds to the value of the $i$th parameter. Moreover, we respectively use $T(\boldsymbol{x})$, $L(\boldsymbol{x})$, $\Phi(\boldsymbol{x})$, and $\Theta(\boldsymbol{x})$ to represent the functions of request throughput, tail latency, average TTFT, and average TPOT for the inference service under configuration $\boldsymbol{x}$. We leverage $\mathcal{C}=\{c_{1},c_{2},...,c_{m}\}$ to symbolize the set of known constraints, where $c_{i}$ is the $i$th known constraint which restricts the relationship between elements of $\boldsymbol{x}$. For example, a known constraint can be expressed as “$\boldsymbol{x}[1]<\boldsymbol{x}[3]\ \text{if}\ \boldsymbol{x}[4]\ \text{is}\ \text{False}$”. Moreover, we utilize $POF(\boldsymbol{x})$ to represent $\boldsymbol{x}$’s probability of feasibility (POF) that $\boldsymbol{x}$ won’t violate hidden constraints. Therefore, the tuning problem can be formulated as
$\mathbb{P}:\max\limits_{\boldsymbol{x}\in\boldsymbol{\Lambda}}\ \ \lambda_{t}\cdot T(\boldsymbol{x}),\lambda_{l}\cdot L(\boldsymbol{x}),\lambda_{\phi}\cdot\Phi(\boldsymbol{x}),\lambda_{\theta}\cdot\Theta(\boldsymbol{x})$ (1)
@@
$c_{i},\forall c_{i}\in\mathcal{C},$ (2)
$POF(\boldsymbol{x})\geq\Delta,$ (3)
@@
IV SCOOT: Solution Description
In this section, we first introduce the overall workflow of SCOOT. Then, we elaborate on how SCOOT employs SOBO and MOBO to address the tuning problem $\mathbb{P}$. Next, we introduce two novel features of SCOOT, the parallel suggestion that can improve the tuning efficiency and random forest-based POF learning that is proposed to learn hidden constraints during the tuning process. Lastly, we describe how SCOOT ensures the stability of optimized SLOs.
IV-A SCOOT Workflow

### v2

# SCOOT: SLO-Oriented Performance Tuning for LLM Inference Engines

@@ Abstract.
As large language models (LLMs) are gaining increasing popularity across a wide range of web applications, it is of great importance to optimize service-level objectives (SLOs) for LLM inference services to enhance user satisfaction and improve the competitiveness of cloud vendors. In this paper, we observe that adjusting the parameters of LLM inference engines can improve service performance, and the optimal parameter configurations of different services are different. Therefore, we propose SCOOT, an automatic performance tuning system to optimize SLOs for each LLM inference service by tuning the parameters of the inference engine. SCOOT jointly exploits single-objective and multiple-objective Bayesian optimization (BO) techniques to handle various optimization objectives via exploration and exploitation. Moreover, SCOOT prunes the search space with known constraints and adopts a random forest to learn hidden constraints during the tuning process to mitigate invalid exploration. To improve the tuning efficiency, SCOOT utilizes the parallel suggestion to accelerate the tuning process. Extensive experiments demonstrate that SCOOT considerably outperforms existing tuning techniques in SLO optimization while greatly improving the tuning efficiency. Moreover, SCOOT is universally applicable to various LLM inference engines including vLLM and TensorRT-LLM. Currently, SCOOT has already been implemented in the production environment at Ant Group.
Service-Level Objective, LLM Inference Engine, Performance Tuning, Bayesian Optimization, Cloud Computing
journalyear: 2025copyright: acmlicensedconference: Proceedings of the ACM Web Conference 2025; April 28–May 2, 2025; Sydney, NSW, Australia.booktitle: Proceedings of the ACM Web Conference 2025 (WWW ’25), April 28–May 2, 2025, Sydney, NSW, Australiaisbn: 979-8-4007-1274-6/25/04doi: 10.1145/3696410.3714930ccs: Computing methodologies Parallel computing methodologiesccs: Computing methodologies Machine learningccs: Computing methodologies Natural language generation
1. Introduction
@@
With the impressive capabilities demonstrated by large language models (LLMs) across various web applications (min2023recent, ), cloud vendors such as Alibaba Cloud and AWS have started offering services of LLM deployment and inference (alibaba_bailian, ; aws_sagemaker, ). Customers (e.g., developers and providers of web applications) can deploy specified LLMs on their exclusive computing resources. Through the application programming interface (API) provided by the cloud vendor, customers’ requests can be continuously served by the deployed LLMs.
@@
These LLM inference services run LLM instances with advanced inference engines such as vLLM (vllm, ) and TensorRT-LLM (tensorrt_llm, ), which are equipped with cutting-edge technologies, such as continuous batching (yu2022orca, ), paged attention (kwon2023efficient, ), and chunked prefill (agrawal2024taming, ). These techniques can accelerate inference and improve throughput, thus delivering a high service performance to customers.
@@
To ensure service performance, customers always agree on service level objectives (SLOs) with cloud vendors. SLOs are defined as a series of performance metric constraints, such as requiring that the 95th percentile latency of requests be less than 1 second. If cloud vendors violate these SLOs, they not only have to compensate customers but also face reputation damage. Therefore, appropriately setting SLOs is critical for cloud vendors (xu2024deploying, ). They should optimize SLOs (e.g., guarantee a lower tail latency) to improve customer satisfaction while ensuring that there won’t be SLO violations.
@@
Typically, cloud vendors stress test services with high request rates and set the performance metrics achieved under the stress testing as SLOs. This ensures that SLOs won’t be violated even under heavy workload scenarios. By improving the service performance under stress testing, cloud vendors can deliver better SLOs to customers. In this paper, we observe that adjusting the parameters of LLM inference engines has great potential to improve service performance, and the optimal parameter configurations of various LLM inference services are different. Therefore, we stand in the shoes of cloud vendors and optimize SLOs for LLM inference services by tuning inference engine parameters under stress testing.
@@
Numerous performance tuning methods have been proposed and applied across various fields (herodotou2020survey, ), but they all fall short of both efficiency and optimality for tuning LLM inference engines. Methods like random sampling and meta-heuristic algorithms, such as Monte Carlo sampling (homem2014monte, ) and genetic algorithms (katoch2021review, ) lack efficiency as they fail to fully utilize historical information. Besides, heuristic searches rely on expert knowledge and elaborate pre-profiling to model relationships between parameters and performance. Nevertheless, with the rapid evolution of technologies in LLM inference engines, the parameters’ numbers and ranges are frequently updated, which makes the modeled parameter-performance relationship outdated and the designed heuristic search methods inapplicable. Additionally, learning-based approaches such as reinforcement learning (RL) (padakandla2021survey, ) and Bayesian optimization (BO) (wang2023recent, ) have also been widely exploited in performance tuning. They can effectively leverage historical information and tune parameters automatically without prior knowledge. However, RL requires a time-consuming training process, while existing BO solutions fail to address the following three challenges.
@@
Challenge 1: Various Optimization Objectives. Customers seek to enhance different performance metrics depending on their application requirements, such as improving request throughput for the periodically invoked offline recommendation application, reducing request tail latency for the online classification application, and minimizing time-to-first-token (TTFT) and time-per-output-token (TPOT) simultaneously for interactive applications such as chatbot, requiring the tuner to have the capability of handling both single-objective and multi-objective optimization problems.
Challenge 2: Complex Known and Hidden Constraints. Some parameters of inference engines depend on the settings of other parameters, thus causing constraints on the search space. For example, for vLLM, max-num-batched-tokens must be greater than or equal to max-num-seqs. We refer to these as known constraints, which can be provided to the tuner ahead of time. Besides, given a specific service, certain parameter combinations can lead to inference engine crashes during stress testing. In this paper, these infeasible parameter combinations are referred to as hidden constraints. Different inference services have different hidden constraints that are initially unknown and must be learned throughout the tuning process. Specifically, for some certain services, vLLM often crashes due to a timeout error during the stress testing when scheduler-delay-factor is set to a large value.
Challenge 3: High Evaluation Overhead. Learning-based tuners always learn the correlation between the service performance and the engine’s parameter settings by evaluating various parameter configurations. As stress testing an inference service takes about 5 to 10 minutes, even with only 30 evaluations for tuning, the total time spent tuning a single service can range from 2.5 to 5 hours. Since the increasing popularity of LLMs leads to the deployment of a large number of LLM inference services, the cumulative time required for tuning these services is prohibitive.
Table 1. vLLM PARAMETERS TO TUNE.
@@
To tackle these challenges, we propose SCOOT, a ServiCe-level Objective Oriented performance Tuning system, which automatically tune parameters of LLM inference engines to optimize SLOs for LLM inference services. We first propose a general formulation of the inference engine tuning problem to accommodate various optimization objectives and complex constraints, and SCOOT can resolve the problem with BO, where single-objective BO (SOBO) and multi-objective BO (MOBO) (khan2002multi, ; yang2019multi, ; daulton2020differentiable, ) are respectively employed to search optimized parameter configurations for single-objective and multi-objective optimization scenarios, thus addressing challenge 1. Since constraint violations result in invalid observations caused by engine crashes, which greatly hurts the tuning efficiency, SCOOT prunes the search space with known constraints and exploits a random forest to learn hidden constraints during the tuning process, thus mitigating challenge 2. To resolve challenge 3, SCOOT employs the parallel suggestion technique to recommend multiple parameter configurations each time for simultaneous evaluation, thus fully utilizing idle computing resources to speed up tuning. SCOOT can support various LLM inference engines including vLLM and TensorRT-LLM. Nowadays, SCOOT is in use at Ant Group.
We conduct extensive experiments with various LLMs and different types and numbers of GPUs under request traces collected from various LLM-based web applications at Ant Group. The results show that SCOOT can speed up the tuning process and significantly optimize SLOs, improving the request throughput by up to 68.3%, reducing the request tail latency by up to 40.6%, and reducing the TTFT and TPOT by up to 99.8% and 61.0%, respectively, compared to the default parameter configuration and existing tuning methods. SCOOT is open-sourced at https://github.com/Ketonmi/SCOOT. The main contribution of this paper is summarized as follows.
  • Random forest regression is employed by SCOOT to learn hidden constraints during the tuning process to avoid invalid explorations, while the parallel suggestion technique is adopted to significantly improve the tuning efficiency using additional computing resources.
2. Background and Motivation
Background: parameters of the LLM inference engine. To provide flexibility of use, inference engines expose many parameters. For vLLM, these parameters include boolean variables such as enable-chunked-prefill that can enable the chunked prefill technique, integer and float variables such as max-num-seqs and scheduler-delay-factor that can change the request scheduling strategy, and enumeration variables such as block-size that can change the memory allocation policy. In this paper, we focus on tuning parameters that do not affect the accuracy of LLMs. Therefore, we do not consider parameters related to model compression, such as quantization. Besides, although speculative decoding has been theoretically proven not to hurt LLM accuracy (leviathan2023fast, ), it still affects the LLMs’ generation results, making it inapplicable for certain applications. Hence, we also do not tune parameters related to speculative decoding. Due to the open-source nature, ease of use, and popularity of vLLM, it is primarily utilized as the inference engine in this paper, and the vLLM parameters to be tuned are listed in Table 1, which constructs a huge search space of over one hundred billion configuration points.
Table 2. VARIOUS SERVICE CHARACTERISTICS.
@@ E BOT A10 4 LLAMA2-7B
Figure 1. Optimal TTFT and TPOT for various services. The TTFT and TPOT shown are relative values compared to those of the default parameter configuration. The lower, the better.
@@
Motivation: parameter adjustment of the LLM inference engine can enhance performance for LLM inference services. We conduct experiments to find the optimal parameter configurations of vLLM with grid searches for five services of various LLMs deployed on different numbers and types of GPUs under request traces of various applications. The service characteristics in the experiments are shown in Table 2, where A10 and A100 separately represent the NVIDIA A10 24G GPU and NVIDIA A100 80G GPU, and SQL and BOT represent the request traces of text-to-SQL and chatbot applications at Ant Group. Other experimental settings are the same as those in Section 5. Figure 1 presents the relative TTFT and TPOT performed by optimal parameter configurations for these services. We can observe that TTFT and TPOT can be reduced by up to 98.9% and 49.9% compared to the default configuration, respectively, which confirms the significance of performance tuning for LLM inference engines.
@@
Motivation: different inference services’ optimal parameter configurations are different. We conduct experiments to apply the optimal parameter configuration of a service to other services. Figure 2 presents the experimental results. We can observe that the performance exhibited by the optimal parameter configuration for a service may perform poorly or even violate hidden constraints and cause errors for other services. Hence, there is no one “best practice” configuration that works best in all scenarios, and it is essential to conduct performance tuning for each inference service.
3. Problem Formulation
@@
Given an inference service, the goal of performance tuning for the LLM inference engine is to find the optimal or near-optimal parameter configuration that maximizes the objective and satisfies known and hidden constraints. As we mentioned before, different customers want to optimize different performance metrics. Therefore, we provide a generalized problem formulation that supports various optimization objectives and complex constraints.
(a) Relative time-to-first-token (TTFT) compared to the default configuration. The lower, the better.
(b) Relative time-per-output-token (TPOT) compared to the default configuration. The lower, the better.
Figure 2. TTFT and TPOT for applying optimal parameter configurations of different services to other services. Conf. $X$ represents the optimal parameter configuration for the service $X\in\{A,B,C,D,E\}$. The red lines (1.0×) in Sub-figures (a) and (b) indicate the TTFT and TPOT under the default parameter configuration of the inference engine, respectively.
@@
Figure 3. SCOOT workflow. SCOOT leverages BO to find optimized parameter configurations via exploration and exploitation.
@@
We define the configuration search space $\boldsymbol{\Lambda}=\Lambda_{1}\times\Lambda_{2}\times...\times\Lambda_{n}$, where $\Lambda_{i}$ represent the range of the $i$th parameter to tune. Vector $\boldsymbol{x}\in\boldsymbol{\Lambda}$ is utilized to denote a specific parameter configuration, where the $i$th element $\boldsymbol{x}[i]\in\Lambda_{i}$ corresponds to the value of the $i$th parameter. Moreover, we respectively use $T(\boldsymbol{x})$, $L(\boldsymbol{x})$, $\Phi(\boldsymbol{x})$, and $\Theta(\boldsymbol{x})$ to represent the functions of request throughput, tail latency, average TTFT, and average TPOT for the inference service under configuration $\boldsymbol{x}$. We leverage $\mathcal{C}=\{c_{1},c_{2},...,c_{m}\}$ to symbolize the set of known constraints, where $c_{i}$ is the $i$th known constraint which restricts the relationship between elements of $\boldsymbol{x}$. For example, a known constraint can be expressed as “$\boldsymbol{x}[1]<\boldsymbol{x}[3]\ \text{if}\ \boldsymbol{x}[4]\ \text{is}\ \text{False}$”. Moreover, we utilize $POF(\boldsymbol{x})$ to represent $\boldsymbol{x}$’s probability of feasibility (POF) that $\boldsymbol{x}$ won’t violate hidden constraints. Therefore, the tuning problem is formulated as follows:
(1) $\mathbb{P}:\max\limits_{\boldsymbol{x}\in\boldsymbol{\Lambda}}\ \ \lambda_{t}\cdot T(\boldsymbol{x}),\lambda_{l}\cdot L(\boldsymbol{x}),\lambda_{\phi}\cdot\Phi(\boldsymbol{x}),\lambda_{\theta}\cdot\Theta(\boldsymbol{x})$
@@
(2) $c_{i},\forall c_{i}\in\mathcal{C},$
(3) $POF(\boldsymbol{x})\geq\Delta,$
@@
4. SCOOT: Solution Description
We present SCOOT’s design, including its use of SOBO and MOBO for solving $\mathbb{P}$, its innovative features of parallel suggestion and random forest-based POF learning, and the SLO robustness assurance.
4.1. SCOOT Workflow